# Dark Matter Freeze-In and Small-Scale Observables: Novel Mass Bounds and Viable Particle Candidates

### v2

@@ I Introduction
As is well known, when DM interactions with the bath are renormalizable, freeze-in production is most efficient when the temperature is comparable to the mass of the heaviest particle involved in the process [6, 7, 8, 9]. This implies that the DM relic abundance can be predicted solely in terms of particle physics parameters such as masses and couplings that could, in principle, be tested in laboratory experiments, without sensitivity to the unknown details of inflationary reheating. This feature is described by saying that freeze-in production is “IR-dominated”. This is undoubtedly a strength of the framework. It is also important to recognize how the feeble interactions render it extremely challenging to test experimentally. Promising avenues for discovery include displaced events observed at particle accelerators and colliders [10, 11, 12, 13, 14, 15, 16, 17, 18, 19], as well as cosmological signatures imprinted on small-scale structure [20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
The main subject of this paper is the latter effect. Freeze-in models can feature large free-streaming lengths, which erase perturbations below a characteristic scale and consequently suppress the formation of small-scale structure. It is therefore essential to employ probes that are sensitive to the DM distribution at these scales. Several observational strategies have been used to constrain DM properties. In particular, observations of ultra-faint dwarf galaxies, the Lyman-$\alpha$ forest, strong gravitational lensing, and stellar streams have all been successfully employed to place limits on models that suppress or enhance structure on small scales [30, 31]. It is important to recognize that these constraints are complementary, as they rely on different observational techniques and are affected by distinct systematic uncertainties. Our work examines the implications of this collection of small-scale probes.
@@ II Warm Dark Matter Mass Bounds
[…]
@@ III Setting mass bounds on quasi-thermal PSD: general strategy
A more comprehensive analysis would employ the full PSD to evaluate the DM impact on the formation of cosmological structures. However, this approach is computationally expensive for deriving bounds in concrete models and impractical for a systematic exploration of their parameter space. Instead, we build on the results of Ref. [25], which established a connection between freeze-in DM and WDM. Using the Boltzmann solver CLASS [76], it was shown in a model-independent manner that the matter power spectrum for freeze-in DM exhibits features closely resembling those of WDM. In particular, the relative “transfer function” [^1] Following standard conventions, we define the squared transfer function as the ratio of the linear power spectrum for freeze-in DM to that of $\Lambda$CDM: $\mathcal{T}^{2}(k)\equiv P(k)/P_{\Lambda\rm CDM}(k)$. displays a sharp cutoff at large wavenumbers, indicating a suppression of small-scale structure that mirrors the behavior of WDM. As a result, the freeze-in transfer function for a given model can be effectively mapped onto that of WDM by selecting an appropriate value of $m_{\rm WDM}$. Our strategy is thus supported qualitatively by the thermal origin of freeze-in DM and quantitatively by the results of Ref. [25].
@@
[…]
First, we note that the quasi-thermal PSD is defined up to an overall normalization fixed by the relic density. Second, the Friedmann-Robertson-Walker (FRW) geometry ensures that the PSD depends only on FRW time and the magnitude of the physical momentum. We express the PSD in terms of the comoving momentum $q$, formally defined in the next paragraph. Importantly, $q$ remains constant once DM production shuts off, naturally factoring out cosmological redshift. Consequently, the functional form in Eq. (1) describes the asymptotic, time-independent PSD in the late universe. The parameters $(\alpha,\beta,\kappa)$ are real, with $\kappa$ restricted to $\{-1,0,+1\}$. For instance, WDM corresponds to a thermal PSD with $(\alpha,\beta)=(0,1)$ and $\kappa$ determined by particle statistics. We require $\beta>0$ for convergence of the normalization integral and $|\alpha|\lesssim 1$ to keep the PSD close to thermal. This is true for all freeze-in scenarios studied in this work.
@@
where $a(T)$ is the FRW scale factor expressed as a function of the bath temperature $T$ via the relation in Eq. (3). The first equality allows us to recast the comoving momentum as $q=p(t)/T_{\chi}$, while the second follows from entropy conservation ($sa^{3}=\text{const}$). As pointed out by Ref. [23] and quantitatively confirmed by Ref. [25], a key quantity for cosmological structure formation is the DM root-mean-square velocity
@@
$m_{\chi}^{\rm min}=22.4\ {\rm keV}\left(\frac{m_{\rm WDM}^{\rm min}}{6\ {\rm keV}}\right)^{4/3}\left(\frac{\sigma_{q}}{3.6}\right)\left(\frac{106.75}{g_{\star s}(T_{P})}\right)^{1/3}.$ (7)
@@ IV Freeze-in production
Bath particles ${\cal B}_{i}$ are described by equilibrium distributions $f^{\rm eq}_{{\cal B}_{i}}(k_{i})=[\exp(E_{i}/T)\pm 1]^{-1}$, either characterized by Fermi-Dirac ($+$) or Bose-Einstein (--) statistics. The PSD evolution is governed by the Boltzmann equation
@@
×$\qquad\qquad\times\prod_{i=1}^{n}f_{{\cal B}_{i}}(k_{i})\prod_{j={n+1}}^{n+m}[1\mp f_{{\cal B}_{j}}(k_{j})]\ .$ (10)
@@ V.2 Sterile Neutrinos
$m_{N}>\left\{\begin{array}[]{ll}\begin{cases}47\,{\rm keV}&\text{MW Satellites}\\ 40\,{\rm keV}&\text{JWST}\\ 37\,{\rm keV}&\text{Lyman}\mathrm{-}\alpha\end{cases}&\quad m_{\rho}=100\,{\rm MeV}\\[30.00005pt] \begin{cases}24\,{\rm keV}&\text{MW Satellites}\\ 21\,{\rm keV}&\text{JWST}\\ 19\,{\rm keV}&\text{Lyman}\mathrm{-}\alpha\end{cases}&\quad m_{\rho}=10\,{\rm GeV}\\[30.00005pt] \begin{cases}22\,{\rm keV}&\text{MW Satellites}\\ 17\,{\rm keV}&\text{JWST}\\ 19\,{\rm keV}&\text{Lyman}\mathrm{-}\alpha\end{cases}&\quad m_{\rho}=1\,{\rm TeV}\end{array}\right.\ .$ (16)
@@
tau $19$ $21$ $24$ $4.9\times 10^{8}$ $7.3\times 10^{-23}$ $1.8\times 10^{36}$
charm $19$ $20$ $23$ $4.5\times 10^{9}$ $1.8\times 10^{-23}$ $3.1\times 10^{37}$
bottom $19$ $21$ $24$ $7.3\times 10^{9}$ $3.2\times 10^{-25}$ $8.3\times 10^{40}$
top $18$ $19$ $22$ $2.8\times 10^{10}$ $1.5\times 10^{-28}$ $5.5\times 10^{47}$
@@ VI.1 Flavor-conserving couplings
For ALP coupled to leptons, $\psi=\ell$, the production channels are lepton/antilepton annihilations $\ell^{+}\ell^{-}\rightarrow\gamma a$ and Compton-like scattering ±$\ell^{\pm}\gamma\rightarrow\ell^{\pm}a$. The squared matrix elements expressed in terms of the Mandelstam variables result in [93, 94, 95]
@@
|±$|{\cal M}_{\ell^{\pm}\gamma\to\ell^{\pm}a}|^{2}=$ $\,\frac{e^{2}}{2\Lambda_{\ell}^{2}}\frac{m_{\ell}^{2}t^{2}}{(s-m_{\ell}^{2})(s+t-m_{\ell}^{2})}\ ,$ (19b)
@@ VIII Conclusions
A simple and efficient way to assess whether small-scale structures are preserved is to compare freeze-in DM with a quasi-thermal PSD to the well-studied case of thermal WDM, which serves as a benchmark for the impact of non-zero free-streaming velocities. In this work, we developed a general methodology based on the criterion introduced in Ref. [23] and numerically validated in Ref. [25], where the key mapping variable is the root-mean-square velocity, $W_{\chi}=\sqrt{\langle{p}\rangle}/m_{\chi}$. This approach allowed us to bypass the need for computationally intensive cosmological Boltzmann solvers such as CLASS [76] when scanning the parameter space of explicit freeze-in models. Thanks to this strategy, we were able to extract lower bounds on the DM mass across a wide range of scenarios.
Figure 7: This figure summarizes all the results obtained in this paper. The variable on the horizontal axis, $\sigma_{q}$, is defined in Eq. (6) as the second moment of the PSD evaluated in terms of the comoving momentum. It quantifies the width of the momentum-space distribution. The vertical axis shows the variable $T_{P}$, which enters the definition of the comoving momentum in Eq. (2). While $T_{P}$ is a conventional parameter, it is convenient to identify it with the temperature at which freeze-in is most efficient—typically corresponding to the mass of the heaviest bath particle participating in DM production. The black isocontours (identical to those in Fig. 3) represent the minimum allowed DM mass produced via freeze-in. These values are obtained through the rescaling given in Eq. (7), using $m_{\rm WDM}^{\rm min}=6.8\,\mathrm{keV}$ as a reference. Each model realization is indicated in the plot by the symbols shown in the legend. In each case, $T_{P}$ is chosen appropriately as the mass of the heaviest particle involved in the decay or scattering process that produces DM. For the dark photon portal, $T_{P}$ is defined by the condition $\omega_{p}(T_{P})=2m_{\chi}$. We show the sterile neutrino model as a dashed cyan line, since the mass of the decaying particle is not specified, therefore $T_{P}$ is also varying; the Higgs portal case is shown by a pink cross. the dark photon portal as a gold star. The dots refer to ALP DM produced by flavor conserving (FC) scatterings of SM fermions with gauge bosons, while the pluses to ALP DM produced by flavor violating (FV) decays of SM fermions. Blue markers indicate processes with the tau lepton, orange with the muon, green with the electron, red with the top quark, violet with the bottom and brown with the charm.
@@
Another important result of this work becomes manifest when comparing the model-independent analysis presented in Fig. 4 with the summary plot shown in Fig. 7. The former was obtained under the assumptions of constant matrix elements for all production processes and Maxwell-Boltzmann statistics for all bath particles. In contrast, for each microscopic realization, we evaluated the transition amplitudes self-consistently and included quantum statistical effects appropriate to the particles involved. Remarkably, the resulting model points align with the model-independent predictions. This agreement is not unexpected: IR ensures that the momentum dependence of the matrix elements has only a minor impact, and Ref. [25] explicitly checked that quantum degeneracy effects can be safely neglected for freeze-in production in the early universe. These findings validate the robustness of our model-independent mass bounds.
To summarize, this work provided an explicit and computationally efficient methodology to derive mass bounds on freeze-in DM in a model-independent manner. The results presented here lay the foundation for future studies involving DM particles produced with a quasi-thermal momentum distribution. Notable examples include scenarios with a subdominant DM component, UV dominated freeze-in production, or freeze-in from a decoupled thermal bath with a temperature different from that of the SM degrees of freedom. We leave these directions for future investigation.
@@ Appendix A The WDM paradigm
$\begin{split}n_{\rm WDM}(t_{0})&=g_{\rm WDM}\int\frac{d^{3}p}{(2\pi)^{3}}f_{\rm WDM}(p)\\ &=g_{\rm WDM}\,g_{\rm eff}\frac{\zeta(3)}{\pi^{2}}T_{\rm WDM}^{3}\,,\end{split}$ (29)

### v3

@@ I Introduction
As is well known, when DM interactions with the bath are renormalizable, freeze-in production is most efficient when the temperature is comparable to the mass of the heaviest particle involved in the process [6, 7, 8, 9]. This implies that the DM relic abundance can be predicted solely in terms of particle physics parameters such as masses and couplings that could, in principle, be tested in laboratory experiments, without sensitivity to the unknown details of inflationary reheating. This feature is described by saying that freeze-in production is “IR-dominated”. This is undoubtedly a strength of the framework. It is also important to recognize how the feeble interactions render it extremely challenging to test experimentally. Promising avenues include displaced events at colliders [10, 11, 12, 13, 14, 15, 16, 17, 18, 19], direct detection [20, 21, 22, 23, 24], and cosmological signatures imprinted on small-scale structure [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40].
The main subject of this paper is the latter effect. Freeze-in models can feature large free-streaming lengths, which erase perturbations below a characteristic scale and consequently suppress the formation of small-scale structure. It is therefore essential to employ probes that are sensitive to the DM distribution at these scales. Several observational strategies have been used to constrain DM properties. In particular, observations of ultra-faint dwarf galaxies, the Lyman-$\alpha$ forest, strong gravitational lensing, and stellar streams have all been successfully employed to place limits on models that suppress or enhance structure on small scales (see, e.g., the reviews provided by Refs. [41, 42] and the references cited in Sec. II). It is important to recognize that these constraints are complementary, as they rely on different observational techniques and are affected by distinct systematic uncertainties. Our work examines the implications of this collection of small-scale probes.
@@ II Warm Dark Matter Mass Bounds
Updated Milky Way satellite counts. The number of satellite galaxies in the MW is expected to be reduced in the case of WDM, as the formation of low-mass subhalos capable of hosting galaxies is suppressed relative to CDM. Therefore, by comparing the predicted number of satellite galaxies in a given WDM model to the observed count ($N_{\rm obs}$), one can place strong constraints on WDM [43, 44, 45, 46, 47, 48, 49]. Uncertainties in these constraints arise, for instance, from the number of satellites associated with the Large Magellanic Cloud, the halo-galaxy connection [41], and the mass of the MW halo, which is estimated to lie in the range (0.5($1.5)\times 10^{12}\,M_{\odot} [50$, 51, 52, 53, 54, 55, 56].
Conservative limits can be obtained by assuming that all subhalos are capable of forming galaxies and by considering a wide range of MW halo masses. By combining data from the Dark Energy Survey (DES) and the Sloan Digital Sky Survey (SDSS) with $N_{\rm obs}=124^{+40}_{-27}$, Ref. [45] finds $m_{\rm WDM}>2.02$ keV (95% CL) after marginalizing over the MW halo mass. Moreover, using updated data from DES and Pan-STARRS1 with $N_{\rm obs}=270$, Ref. [43] obtains conservative bounds of $m_{\rm WDM}>3.6$$5.1$ keV (95% CL), which vary only weakly with the MW halo mass. After correcting for anisotropic satellite distributions, $N_{\rm obs}$ was updated to $220\pm 50$ [57], which has only a minor impact on the conservative bounds. Incorporating a galaxy formation model leads to significantly stronger constraints, as it further reduces the total number of satellites. For instance, Ref. [44] derives a bound of $m_{\rm WDM}>6.5$ keV (95% CL) using an empirical galaxy–halo model and updated DES and Pan-STARRS1 data with $N_{\rm obs}=220\pm 50$.
@@ III Setting mass bounds on quasi-thermal PSD: general strategy
[…]
@@
$\text{Quasi-thermal PSD:}\qquad f_{\chi}^{\rm qth}(q)\propto\frac{q^{\alpha}}{e^{\beta q}+\kappa}\ .$ (1)
First, we note that the quasi-thermal PSD is defined up to an overall normalization fixed by the relic density. Second, the Friedmann-Robertson-Walker (FRW) geometry ensures that the PSD depends only on FRW time and the magnitude of the physical momentum. We express the PSD in terms of the comoving momentum $q$, formally defined in the next paragraph. Importantly, $q$ remains constant once DM production shuts off, naturally factoring out cosmological redshift. Consequently, the functional form in Eq. (1) describes the asymptotic, time-independent PSD in the late universe. The parameters $(\alpha,\beta,\kappa)$ are real, with $\kappa$ restricted to $\{-1,0,+1\}$. For instance, WDM corresponds to a thermal PSD with $(\alpha,\beta)=(0,1)$ and $\kappa$ determined by particle statistics. We require $\beta>0$ for convergence of the normalization integral and $|\alpha|\lesssim 1$ to keep the PSD close to thermal. This is true for all freeze-in scenarios studied in this work. We stress once again that Eq. (1) is only for illustrative purposes and no results of this work depend on it.
@@
where $a(T)$ is the FRW scale factor expressed as a function of the bath temperature $T$ via the relation in Eq. (3). The first equality allows us to recast the comoving momentum as $q=p(t)/T_{\chi}$, while the second follows from entropy conservation ($sa^{3}=\text{const}$). As pointed out by Refs. [25, 26, 27, 28, 31] and quantitatively confirmed by Ref. [34], a key quantity for cosmological structure formation is the DM root-mean-square velocity
@@
$m_{\chi}^{\rm min}=22\ {\rm keV}\left(\frac{m_{\rm WDM}^{\rm min}}{6\ {\rm keV}}\right)^{4/3}\left(\frac{\sigma_{q}}{3.6}\right)\left(\frac{106.75}{g_{\star s}(T_{P})}\right)^{1/3}.$ (7)
@@ IV Freeze-in production
Bath particles ${\cal B}_{i}$ are described by equilibrium distributions $f^{\rm eq}_{{\cal B}_{i}}(k_{i})=[\exp(E_{i}/T)\pm 1]^{-1}$, either characterized by Fermi-Dirac (++) or Bose-Einstein (-) statistics. The PSD evolution is governed by the Boltzmann equation
@@
$\qquad\qquad\times\prod_{i=1}^{n}f_{{\cal B}_{i}}(k_{i})\prod_{j={n+1}}^{n+m}[1\mp f_{{\cal B}_{j}}(k_{j})]\ .$ (10)
@@ V.2 Sterile Neutrinos
$m_{N}>\left\{\begin{array}[]{ll}\begin{cases}47\,{\rm keV}&\text{MW Satellites}\\ 40\,{\rm keV}&\text{JWST}\\ 37\,{\rm keV}&\text{Lyman}\mathrm{-}\alpha\end{cases}&\quad m_{\rho}=100\,{\rm MeV}\\[30.00005pt] \begin{cases}24\,{\rm keV}&\text{MW Satellites}\\ 21\,{\rm keV}&\text{JWST}\\ 19\,{\rm keV}&\text{Lyman}\mathrm{-}\alpha\end{cases}&\quad m_{\rho}=10\,{\rm GeV}\\[30.00005pt] \begin{cases}22\,{\rm keV}&\text{MW Satellites}\\ 19\,{\rm keV}&\text{JWST}\\ 17\,{\rm keV}&\text{Lyman}\mathrm{-}\alpha\end{cases}&\quad m_{\rho}=1\,{\rm TeV}\end{array}\right.\$. (16)
@@
tau 1919 2121 $24$ $4.9\times 10^{8}$ $7.3\times 10^{-23}$ $1.8\times 10^{36}$
charm 1919 2020 $23$ $4.5\times 10^{9}$ $1.8\times 10^{-23}$ $3.1\times 10^{37}$
bottom 1919 2121 $24$ $7.3\times 10^{9}$ $3.2\times 10^{-25}$ $8.3\times 10^{40}$
top $18$ 1919 $22$ $2.8\times 10^{10}$ $1.5\times 10^{-28}$ $5.5\times 10^{47}$
@@ VI.1 Flavor-conserving couplings
For ALP coupled to leptons, $\psi=\ell$, the production channels are lepton/antilepton annihilations $\ell^{+}\ell^{-}\rightarrow\gamma a$ and Compton-like scattering $\ell^{\pm}\gamma\rightarrow\ell^{\pm}a$. The squared matrix elements expressed in terms of the Mandelstam variables result in [104, 105, 106]
@@
$|{\cal M}_{\ell^{\pm}\gamma\to\ell^{\pm}a}|^{2}=$ $\,\frac{e^{2}}{2\Lambda_{\ell}^{2}}\frac{m_{\ell}^{2}t^{2}}{(s-m_{\ell}^{2})(s+t-m_{\ell}^{2})}\ ,$ (19b)
@@ VII Dark Photon Portal
We conclude this section with a comparison from the analysis of Ref. [32] on the same model. Our bounds are roughly a factor of two more stringent, and we identify this discrepancy as due to different implementations of the criterion used to determine the minimum value of $m_{\chi}$ consistent with WDM clustering bounds. In Ref. [32], $m_{\chi}$ is chosen such that the half-mode (defined as the scale where the DM power spectrum is suppressed by 50% relative to CDM) matches that of the WDM spectrum. This methodology is supported by what highlighted by Ref. [120]. However, as evident in Fig. 3 of Ref. [32], the resulting allowed power spectrum for the non-thermal distribution contains a high-power region that lies below the corresponding WDM spectrum and therefore suppresses power on scales larger than the half-mode. Instead, our criterion for imposing the mass bound on quasi-thermal PSDs follows the criterion established in Ref. [99]. This criterion is further supported by Ref. [34], which shows near-perfect agreement between their area-based criterion, a more sophisticated integrated power criterion, and the warmness bound derived from the second moment $\sigma_{q}$ of the DM PSD. The origin of the discrepancy between these two mass bounds is thus explained: achieving an allowed power spectrum requires a larger high-$k$ cutoff, which in turn implies a higher $m_{\chi}^{\rm min}$. [^3] We have explicitly verified that our bound is consistent with results obtained by inputting the PSD into the Boltzmann solver CLASS [87] and computing the power suppression at small scales.
@@ VIII Conclusions
Figure 7: This figure summarizes all the results obtained in this paper. The variable on the horizontal axis, $\sigma_{q}$, is defined in Eq. (6) as the second moment of the PSD evaluated in terms of the comoving momentum. It quantifies the width of the momentum-space distribution. The vertical axis shows the variable $T_{P}$, which enters the definition of the comoving momentum in Eq. (2). While $T_{P}$ is a conventional parameter, it is convenient to identify it with the temperature at which freeze-in is most efficient—typically corresponding to the mass of the heaviest bath particle participating in DM production. The black isocontours (identical to those in Fig. 3) represent the minimum allowed DM mass produced via freeze-in. These values are obtained through the rescaling given in Eq. (7), using $m_{\rm WDM}^{\rm min}=6.8\,\mathrm{keV}$ as a reference. Each model realization is indicated in the plot by the symbols shown in the legend. In each case, $T_{P}$ is chosen appropriately as the mass of the heaviest particle involved in the decay or scattering process that produces DM. For the dark photon portal, $T_{P}$ is defined by the condition $\omega_{p}(T_{P})=2m_{\chi}$. For the ALP DM scatterings we show the bounds obtained by assuming constant matrix elements as empty circles. We show the sterile neutrino model as a dashed cyan line, since the mass of the decaying particle is not specified, therefore $T_{P}$ is also varying; the Higgs portal case is shown by a pink cross. the dark photon portal as a gold star. The dots refer to ALP DM produced by flavor conserving (FC) scatterings of SM fermions with gauge bosons, while the pluses to ALP DM produced by flavor violating (FV) decays of SM fermions. Blue markers indicate processes with the tau lepton, orange with the muon, green with the electron, red with the top quark, violet with the bottom and brown with the charm.
A simple and efficient way to assess whether small-scale structures are preserved is to compare freeze-in DM with a quasi-thermal PSD to the well-studied case of thermal WDM, which serves as a benchmark for the impact of non-zero free-streaming velocities. In this work, we developed a general methodology based on the criterion introduced in Ref. [31] and numerically validated in Ref. [34], where the key mapping variable is the root-mean-square velocity, $W_{\chi}=\sqrt{\langle{p}^{2}\rangle}/m_{\chi}$. This approach allowed us to bypass the need for computationally intensive cosmological Boltzmann solvers such as CLASS [87] when scanning the parameter space of explicit freeze-in models. Thanks to this strategy, we were able to extract lower bounds on the DM mass across a wide range of scenarios.
@@
Another important result of this work becomes manifest when comparing the model-independent analysis presented in Fig. 4 with the summary plot shown in Fig. 7. The former was obtained under the assumptions of constant matrix elements for all production processes and Maxwell-Boltzmann statistics for all bath particles. In contrast, for each microscopic realization, we evaluated the transition amplitudes self-consistently and included quantum statistical effects appropriate to the particles involved. For the ALP DM produced via binary scatterings, we show the bounds on the DM mass obtained under the constant matrix element— that has been employed in setting model-independent bounds in Sec. IV— as empty circles. Remarkably, the resulting model points align with the model-independent predictions up to 5%-10% accuracy, with the constant matrix element assumption providing a more conservative bound. An important exception is the dark photon portal, for which the assumption of a single production mechanism does not hold due to contamination from plasmon decays. This agreement is not unexpected: IR ensures that the momentum dependence of the matrix elements has only a minor impact, and Ref. [34] explicitly checked that quantum degeneracy effects can be safely neglected for freeze-in production in the early universe. These findings validate the robustness of our model-independent mass bounds under the assumption of no interferences between production mechanisms.
To summarize, this work provided an explicit and computationally efficient methodology to derive mass bounds on freeze-in DM in a model-independent manner. The results presented here lay the foundation for future studies involving DM particles produced with a quasi-thermal momentum distribution. Notable examples include scenarios with a subdominant DM component [34, 121], UV-dominated freeze-in production [122, 26, 123, 124, 33, 125, 126, 127, 128, 129, 130, 131], or freeze-in from a decoupled thermal bath with a temperature different from that of the SM degrees of freedom [132]. We leave these directions for future investigation.
@@ Appendix A The WDM paradigm
where $\zeta(3)$ is the Riemann zeta function. The statistical factor $g_{\rm eff}=\{1,\ 3/4,\ 1/\zeta(3)\}$ arises from the momentum integral of the PSD in the BE, FD, and MB cases, respectively, as listed in Eq. (28).